A mixture of 30 lbs. of candy sells for $1.10 a pound. The mixture consists of chocolates worth $1.50 a pound and chocolates worth 90¢ a pound. How many pounds of each kind were used to make the mixture?
Answer by fanism(39)

let x be the chocolates worth $1.50 a pound
and let y be the chocolates worth 90¢ a pound.
1.5x + 0.9y = 30(1.1) -- eq1
x + y = 30 --> x = 30 - y -- eq2
sub eq2 into eq1, u get
1.5(30-y) + 0.9y = 33
12 = 0.6y
y = 20
put back into eq2,
x + 20 = 30
x = 10